# Integrating Paralinguistics in Speech-Empowered Large Language Models for Natural Conversation

### v2

@@ Abstract
Recent work shows promising results in expanding the capabilities of large language models (LLM) to directly understand and synthesize speech. However, an LLM-based strategy for modeling spoken dialogs remains elusive, calling for further investigation. This paper introduces an extensive speech-text LLM framework, the Unified Spoken Dialog Model (USDM), designed to generate coherent spoken responses with naturally occurring prosodic features relevant to the given input speech without relying on explicit automatic speech recognition (ASR) or text-to-speech (TTS) systems. We have verified the inclusion of prosody in speech tokens that predominantly contain semantic information and have used this foundation to construct a prosody-infused speech-text model. Additionally, we propose a generalized speech-text pretraining scheme that enhances the capture of cross-modal semantics. To construct USDM, we fine-tune our speech-text model on spoken dialog data using a multi-step spoken dialog template that stimulates the chain-of-reasoning capabilities exhibited by the underlying LLM. Automatic and human evaluations on the DailyTalk dataset demonstrate that our approach effectively generates natural-sounding spoken responses, surpassing previous and cascaded baselines. We will make our code and checkpoints publicly available.
@@ 1 Introduction
Large language models (LLMs) have gained significant traction thanks to emergent capabilities [1, 2, 3, 4], such as few-shot in-context learning, complex reasoning [5, 6], and instruction-following [7]. These remarkable discoveries led to chat-enabled LLMs and generative personal assistants [8]. However, text-based agents are limited in usability due to their medium of interaction. Ideally, speech-enabled LLMs would recognize the user’s emotional state or subtle nuance and generate spoken responses with prosody most appropriate to the user’s context. Although automatic speech recognition (ASR) and text-to-speech (TTS) systems can be easily employed, the linguistic discrepancy between speech and text causes dialog inefficiencies and result in sub-optimal user experience [9, 10]. As such, systematically integrating the speech modality into LLMs can unlock speech interactivity while retaining LLMs’ powerful capabilities.
@@ 3.1 Speech-to-Unit Encoder
To further investigate the paralinguistic information contained in the units, we train a separate unit-to-speech reconstruction model using 54,000 hours of speech data. The reconstruction model is trained using the architecture of Voicebox, one of the zero-shot stochastic TTS models [65]. Unlike Voicebox, which takes text and reference speech as inputs for adaptation, our model is trained to generate speech solely from a unit sequence without any reference speech. The comparison between the original audio and the speech reconstructed from the extracted units shows that while the timbre and absolute pitch of the reconstructed speech differ, the pitch variation has a similar trend. This implies a crucial role in conveying non-verbal characteristics such as emotions, which closely match the original audio, as shown in Figure 2. Additionally, we have uploaded samples of several ground truth audios and corresponding reconstructed audios on our demo page.
Through these two experiments, we confirm that the acoustic units, commonly known to primarily encode semantic information, also contain a significant amount of paralinguistic information, such as emotions and pitch variations. We adopt this speech tokenization scheme for our speech-text pretraining and spoken dialog fine-tuning to help capture non-verbal cues in spoken conversations. More detailed descriptions of these experiments are provided in Appendix A.3.1 and A.3.4.
@@ 3.2 Unified Speech-Text Pretraining
In this section, we introduce a unified speech-text pretraining scheme that extends pretrained LLM to speech-text cross-modality. Our overall speech-text pretraining scheme is in Figure 4 in the Appendix.
@@ 5 Conclusion
Despite these advantages, our model has several limitations and areas for improvement. Firstly, the exploration of datasets and models used for pretraining is limited. Further investigation is necessary to determine which data are crucial for our pretraining scheme and to explore whether our pretraining scheme could be effective with other LLMs beyond Mistral-7B. Secondly, building a spoken dialog model capable of directly generating spoken responses from input spoken dialog without the need for cross-modal chaining can be a promising direction. Lastly, we also plan to investigate whether our pretraining approach is beneficial for other speech-text tasks beyond spoken dialog modeling.
@@ A.3.1 Emotional Cues in Acoustic Units
A.3.2 Overview of Unified Speech-Text Pretraining
We have illustrated an overview of the proposed unified speech-text pretraining in Figure 4.
Figure 4: The overall speech-text pretraining scheme.
A.3.3 Templates for Fine-tuning
@@
Figure 5 illustrates the template of single-turn spoken dialog and text dialog we used for fine-tuning. We perform loss calculations only on the highlighted part.

### v3

# Paralinguistics-Aware Speech-Empowered Large Language Models for Natural Conversation

@@ Abstract
Recent work shows promising results in expanding the capabilities of large language models (LLM) to directly understand and synthesize speech. However, an LLM-based strategy for modeling spoken dialogs remains elusive, calling for further investigation. This paper introduces an extensive speech-text LLM framework, the Unified Spoken Dialog Model (USDM), designed to generate coherent spoken responses with naturally occurring prosodic features relevant to the given input speech without relying on explicit automatic speech recognition (ASR) or text-to-speech (TTS) systems. We have verified the inclusion of prosody in speech tokens that predominantly contain semantic information and have used this foundation to construct a prosody-infused speech-text model. Additionally, we propose a generalized speech-text pretraining scheme that enhances the capture of cross-modal semantics. To construct USDM, we fine-tune our speech-text model on spoken dialog data using a multi-step spoken dialog template that stimulates the chain-of-reasoning capabilities exhibited by the underlying LLM. Automatic and human evaluations on the DailyTalk dataset demonstrate that our approach effectively generates natural-sounding spoken responses, surpassing previous and cascaded baselines. Our code and checkpoints are available at https://github.com/naver-ai/usdm.
@@ 1 Introduction
Large language models (LLMs) have gained significant traction thanks to emergent capabilities [1, 2, 3, 4, 5], such as few-shot in-context learning, complex reasoning [6, 7], and instruction-following [8]. These remarkable discoveries led to chat-enabled LLMs and generative personal assistants [9]. However, text-based agents are limited in usability due to their medium of interaction. Ideally, speech-enabled LLMs would recognize the user’s emotional state or subtle nuance and generate spoken responses with prosody most appropriate to the user’s context. Although automatic speech recognition (ASR) and text-to-speech (TTS) systems can be easily employed, the linguistic discrepancy between speech and text causes dialog inefficiencies and result in sub-optimal user experience [10, 11]. As such, systematically integrating the speech modality into LLMs can unlock speech interactivity while retaining LLMs’ powerful capabilities.
@@ 3.1 Speech-to-Unit Encoder
To further investigate the paralinguistic information contained in the units, we train a separate unit-to-speech reconstruction model using 54,000 hours of speech data. The reconstruction model is trained using the architecture of Voicebox, one of the zero-shot stochastic TTS models [66]. Unlike Voicebox, which takes text and reference speech as inputs for adaptation, our model is trained to generate speech solely from a unit sequence without any reference speech. The comparison between the original audio and the speech reconstructed from the extracted units shows that while the timbre and absolute pitch of the reconstructed speech differ, the pitch variation has a similar trend, as shown in Figure 2. This implies a crucial role in conveying non-verbal characteristics such as emotions, which closely match the original audio. Additionally, we have uploaded samples of several ground truth audios and corresponding reconstructed audios on our demo page.
Through these two experiments, we confirm that the acoustic units, commonly known to primarily encode semantic information, also contain a significant amount of paralinguistic information, such as emotions and pitch variations. We adopt this speech tokenization scheme for our speech-text pretraining and spoken dialog fine-tuning to help capture non-verbal cues in spoken conversations. More detailed descriptions of these experiments are provided in Appendix A.3.1 and A.3.3.
@@ 3.2 Unified Speech-Text Pretraining
Figure 3: The overall speech-text pretraining scheme.
In this section, we introduce a unified speech-text pretraining scheme that extends pretrained LLM to speech-text cross-modality. Our overall speech-text pretraining scheme is in Figure 3.
@@ 5 Conclusion
Despite these advantages, our model has several limitations and areas for improvement. Firstly, the exploration of datasets and models used for pretraining is limited. Further investigation is necessary to determine which data are crucial for our pretraining scheme and to explore whether our pretraining scheme could be effective with other LLMs beyond Mistral-7B. Secondly, building a spoken dialog model capable of directly generating spoken responses from input spoken dialog without the need for cross-modal chaining can be a promising direction. Next, the current pretraining scheme is based on tens of thousands of hours of English data, and it has limitations when applied to other languages with relatively smaller amounts of speech data compared to English. We plan to expand our model to a variety of languages in addition to English. Lastly, we also plan to investigate whether our pretraining approach is beneficial for other speech-text tasks beyond spoken dialog modeling.
Acknowledgments and Disclosure of Funding
This work was supported by SNU-Naver Hyperscale AI Center, Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720, No.2022R1A5A7083908), and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, SNU in 2024.
@@ A.3.1 Emotional Cues in Acoustic Units
A.3.2 Templates for Fine-tuning
@@
Figure 5 illustrates the template of single-turn spoken dialog and text dialog we used for fine-tuning. We perform loss calculations only on the highlighted part. USDM refers to our model, and From Scratch and Cascaded are the baselines we used as described in Section 4.1.1.